# Kronecker factorization theorems for the exceptional Malcev algebra

### v2

@@ 4 Some equivalence of categories
Let $\mathcal{M}$ be an arbitrary Malcev algebra and let $f:\mathcal{M}\times\mathcal{M}\longrightarrow F$ denote the non-singular symmetric bilinear form defined in $\mathcal{M}$. As in the previous paragraphs, consider the associative algebra with involution $\mathcal{L}(\mathcal{M})$ with
@@
Let $\mathcal{M}$ be a Malcev algebra with J-admissible involution (that is, every symmetric element lies in the nucleus of $\mathcal{M}$) such that $\mathcal{M}$ contains $\mathbb{M}$ as a subalgebra with $m\mathbb{M}\neq 0$ for any $m\neq 0$ from $\mathcal{M}$. If $f:\mathcal{M}\times\mathcal{M}\longrightarrow F$ is a non-singular symmetric bilinear form defined in $\mathcal{M}$. Then $\mathcal{M}\cong\mathbb{M}\otimes_{F}\mathcal{U}$ for a certain commutative associative algebra $\mathcal{U}$ of symmetric elements.
@@
Now let $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ be an arbitrary Malcev superalgebra and let $h:\mathcal{M}\times\mathcal{M}\longrightarrow F$ denote the non-singular symmetric bilinear form defined in $\mathcal{M}$. Consider the associative superalgebra with superinvolution $\mathfrak{L}(\mathcal{M})$ with
@@
Let $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ be a Malcev superalgebra with J-admissible superinvolution (that is, every symmetric element lies in the nucleus of $\mathcal{M}$) such that $\mathcal{M}_{0}$ contains $\mathbb{M}$, with $m\mathbb{M}\neq 0$ for any $m\neq 0$ from $\mathcal{M}_{0}\cup\mathcal{M}_{1}$. If $h:\mathcal{M}\times\mathcal{M}\longrightarrow F$ is a non-singular symmetric bilinear form defined in $\mathcal{M}$. Then $\mathcal{M}\cong\mathbb{M}\otimes_{F}U$ for a certain supercommutative and associative superalgebra $U$ of symmetric elements.
@@
Consider the superalgebra $\mathfrak{L}(\mathcal{M})$ with superinvolution $\ast$. We can associate the superinvolution $a\longmapsto\overline{a}$ of $\mathcal{M}$ with the superinvolution of $\mathfrak{L}(\mathcal{M})$ by

### v3

@@ 4 Some equivalence of categories
Let $\mathcal{M}$ be a Malcev algebra in $\mathcal{H}$ and let $f:\mathcal{M}\times\mathcal{M}\longrightarrow F$ denote the non-singular symmetric bilinear form defined in $\mathcal{M}$. As in the previous paragraphs, consider the associative algebra with involution $\mathcal{L}(\mathcal{M})$ with
@@
Let $\mathcal{M}$ be a Malcev algebra in $\mathcal{H}$ with J-admissible involution (that is, every symmetric element lies in the nucleus of $\mathcal{M}$) such that $\mathcal{M}$ contains $\mathbb{M}$ as a subalgebra with $m\mathbb{M}\neq 0$ for any $m\neq 0$ from $\mathcal{M}$. If $f:\mathcal{M}\times\mathcal{M}\longrightarrow F$ is a non-singular symmetric bilinear form defined in $\mathcal{M}$. Then $\mathcal{M}\cong\mathbb{M}\otimes_{F}\mathcal{U}$ for a certain commutative associative algebra $\mathcal{U}$ of symmetric elements.
@@
Now let $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ be a Malcev superalgebra in $\mathcal{H}$ and let $h:\mathcal{M}\times\mathcal{M}\longrightarrow F$ denote the non-singular symmetric bilinear form defined in $\mathcal{M}$. Consider the associative superalgebra with superinvolution $\mathfrak{L}(\mathcal{M})$ with
@@
Let $\mathcal{M}=\mathcal{M}_{0}\oplus\mathcal{M}_{1}$ be a Malcev superalgebra in $\mathcal{H}$ with J-admissible superinvolution (that is, every symmetric element lies in the nucleus of $\mathcal{M}$) such that $\mathcal{M}_{0}$ contains $\mathbb{M}$, with $m\mathbb{M}\neq 0$ for any $m\neq 0$ from $\mathcal{M}_{0}\cup\mathcal{M}_{1}$. If $h:\mathcal{M}\times\mathcal{M}\longrightarrow F$ is a non-singular symmetric bilinear form defined in $\mathcal{M}$. Then $\mathcal{M}\cong\mathbb{M}\otimes_{F}U$ for a certain supercommutative and associative superalgebra $U$ of symmetric elements.
@@
Consider the superalgebra $\mathfrak{L}(\mathcal{M})$ with superinvolution $\ast$. We can associate the superinvolution $a\mapsto\overline{a}$ of $\mathcal{M}$ with the superinvolution of $\mathfrak{L}(\mathcal{M})$ by